# SF-Loc: A Visual Mapping and Geo-Localization System based on Sparse Visual Structure Frames

### v1

@@ V-B Fine Pose Estimation
After successful retrieval of the map frame corresponding to the query image, fine-grained re-localization can be performed. In our system, this process could be achieved only based on one single map frame and is independent of any re-constructed model, considering that the visual structure frame contains the full-view depth. This makes the re-localization simple and highly flexible.
@@
Despite the simplicity, the above methods couldn’t be directly used for multi-frame matching. In the application scenarios that our system concerns (cross-season, dynamic objects, etc.), robust feature matching is hard to promise, even with state-of-the-art learning-based techniques. To maximize the success rate and provide accurate pose estimation under large amounts of outliers, we combine multi-frame correspondences and the odometry in a factor graph, as shown in Fig. 6. For better efficiency, we don’t do exhaustive matching among all the related frames. Instead, we compute the correspondences between every query-to-database image pair, based on the VPR results. The cost function of the multi-frame pose estimation is written as follows
$\min\!\sum_{k\in\mathcal{K}}\!\left\|{\mathbf{T}^{w}_{c_{k+1}}}^{\!\!\!\!-1}\!\circ\!\mathbf{T}^{w}_{c_{k}}\!\circ\!\hat{\mathbf{T}}^{c_{k}}_{c_{k+1}}\right\|\!+\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\left\|\lambda^{m_{i}}_{f}\!-\!\hat{\lambda}^{m_{i}}_{f}\!\!\right\|$ (19)
$+\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\left\|\pi_{c}\left(\mathbf{T}^{c_{k}}_{m_{i}}\circ{\pi}_{m}^{-1}(\mathbf{u}^{m_{i}}_{f},\lambda^{m_{i}}_{f})\right)-\mathbf{u}^{c_{k}}_{f}\right\|$ (20)
where $\hat{\mathbf{T}}^{c_{k}}_{c_{k+1}}$ is the relative pose estimation, $\hat{\lambda}^{m_{i}}_{f}$ is the sampled inverse depth of point $f$ on the map frame $m_{i}$, $\mathcal{F}_{k,i}$ is the matched feature point set between map frame $m_{i}$ and user-end frame $c_{k}$, $\mathbf{u}^{m_{i}}_{f}$ and $\mathbf{u}^{c_{k}}_{f}$ are the point coordinates respectively. Notice that the map frame poses are fixed due to unobservability of the reference.
Figure 7: Illustration of how depth uncertainty of the map frame affects the user pose estimation accuracy. In the figure, $\delta s$ denotes the scale uncertainty of the depth map, $\delta p$ denotes the position error of user-end localization. An approximate $\delta p\propto\delta s\cdot|\mathbf{t}^{m_{i}}_{c_{k}}|$ relationship can be found.
@@ VI Experiments
Figure 8: Vehicle trajectories and example images of the experimental dataset. The data sequence collected in 2022.10.23 (5400 sec duration, 30 km mileage, 16 km one-way road coverage) is used for mapping, while the data sequence collected in 2023.04.12 (1800 sec duration, 10 km mileage, 4.9 km one-way road coverage) is used for localization.
@@ VI-A Mapping Performance
For large-scale mapping, obtaining precise global positioning is fundamental. Nevertheless, it is challenging considering the intermittent degradation of GNSS observations in the experimental scenario. In Fig. 9, we show the pose estimation results of different approaches, including: 1) GNSS RTK, 2) VINS-Fusion (monocular VIO + RTK mode, final optimized trajectory) and 3) the proposed multi-sensor DBA scheme in real-time mode and global mode respectively. As can be found, the GNSS RTK results are noisy. Although decimeter-level accuracy is intermittently met, gross errors occur frequently. By integrating RTK with VIO, the VINS-Fusion scheme suppresses the gross errors to some extent. For the proposed DBA (real-time) scheme, by integrating visual DBA, inertial information and RTK in a tightly-coupled way, the GNSS errors are better mitigated. Furthermore, with the global optimization presented in Sect. IV-A, the position estimation error the DBA (global) scheme is mostly kept to below 1 m throughout the sequence, which ensures the decimeter-level lower bound of user-end absolute localization.
@@
    3 37.67 MB (JPEG) + 3.41 MB (Global Desc.) + 7.76 MB (Depth)
    4 25.81 MB (JPEG) + 2.32 MB (Global Desc.) + 5.30 MB (Depth)
@@
On the other hand, the generation of map data is looked into. As baselines, we test ORBSLAM3 and HLoc for the mapping task, which are both commonly used pipelines for visual mapping and re-localization. For ORBSLAM3, we simply use the monocular visual-inertial mode to process the sequence and save the map data in the Atlas format. The map storage is over 6 GB for the whole sequence, which is relatively heavy for practical applications. As for HLoc, we found it hard to achieve such large-scale SfM through the default COLMAP + SIFT pipeline. Instead, we directly use the ground-truth camera poses and perform feature extraction, matching and triangulation based on Superpoint and SuperGlue. The map storage is over 30 GB, which is 5 times larger than the ORBSLAM3 map. Such heavy map makes it hard to be applied to large-scale scenarios.
@@ VI-B Localization Performance
As to user-end localization, we check the coarse-to-fine localization performance of the proposed system.
@@ VI-B1 Coarse Localization
We also list the results based on a more sparse map with co-visibility threshold $\Xi=0.3$. Generally, the recall is noticeably lower than $\Xi=0.3$ condition, while the EigenPlaces50 (SSS10) scheme can still achieve 93.8%/100% 10m/20m recall.
@@ VI-B2 Fine Localization
As to the evaluation of fine localization, we test the re-localization modes of ORB-SLAM3 and HLoc as the baselines, and also test GNSS-based schemes for reference, including single point positioning (SPP) and VINS-Fusion (stereo VIO + SPP). For the proposed SF-Loc, we test both single-frame and multi-frame re-localization modes. For multi-frame re-localization, the user-end relative poses are provided by stereo VIO. The factor graph optimization implementation in Sect. V-B is used; For comparison, we also test PnP with RANSAC upon the SF-Loc map.
For ORB-SLAM3 and HLoc, the coarse-to-fine localization is performed based on their own pipeline and map data (corresponding to TABLE I). For SF-Loc, the map data generated with different co-visibility thresholds (0.3 and 0.4) are considered, and the coarse localization results of EigenPlaces50 (SSS10) are used.
@@
Finally, we show the time consumption statistics of user-end localization in TABLE IV, which is tested on a laptop with Nvidia RTX 4080 GPU.
TABLE IV: Time Cost of Different Modules in User-End Localization.
@@ VII Conclusion
In this paper, we propose a vision-centered mapping and localization framework based on the map representation of visual structure frames. The highlight of the system is the lightweight map storage by making the frame data compact and keeping the sparsity, while can still support high-recall, high-accuracy localization through the utilization of user-end multi-frame information.

### v2

@@ V-B Fine Pose Estimation
After successful retrieval of the map frame corresponding to the query image, fine-grained re-localization can be performed. In our system, this process could be achieved based on one single map frame and is independent of any re-constructed model, considering that the visual structure frame contains the full-view depth. This makes the re-localization simple and highly flexible.
@@
Despite the simplicity, the above methods couldn’t be directly used for multi-frame matching. In the application scenarios that our system concerns (cross-season, dynamic objects, etc.), robust feature matching is hard to promise, even with state-of-the-art learning-based techniques. To maximize the success rate and provide accurate pose estimation under large amounts of outliers, we combine multi-frame correspondences and the odometry in a factor graph, as shown in Fig. 6. For better efficiency, we don’t do exhaustive matching among all the related frames. Instead, we compute the correspondences between every query-to-database image pair from VPR. Based on the re-projection errors, the cost function of the multi-frame pose estimation is written as follows
$\begin{split}\min\!\sum_{k\in\mathcal{K}}\!\!\left\|{\mathbf{T}^{w}_{c_{k+1}}}^{\!\!\!\!\!\!-1}\!\!\circ\!\mathbf{T}^{w}_{c_{k}}\!\!\circ\!\hat{\mathbf{T}}^{c_{k}}_{c_{k+1}}\!\right\|^{2}_{\bm{\Omega}_{p}}\!\!\!\!+\!\!\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\!\left\|\lambda^{m_{i}}_{f}\!\!-\!\!\hat{\lambda}^{m_{i}}_{f}\!\right\|^{2}_{\Omega_{\lambda}}\\ \!\!\!+\!\!\sum_{k\in\mathcal{K}}\!\sum_{f\in\mathcal{F}_{k,i}}\!\!\!\rho_{C}\!\!\left(\left\|\pi_{c}\!\left(\!\mathbf{T}^{c_{k}}_{m_{i}}\!\!\circ\!{\pi}_{m}^{-1}(\mathbf{u}^{m_{i}}_{f},\lambda^{m_{i}}_{f})\right)\!-\!\mathbf{u}^{c_{k}}_{f}\right\|^{2}_{\bm{\Omega}_{c}}\right)\end{split}$ (19)
where $\hat{\mathbf{T}}^{c_{k}}_{c_{k\!+\!1}}$ is the relative pose estimation, $\hat{\lambda}^{m_{i}}_{f}$ is the sampled inverse depth of point $f$ on the map frame $m_{i}$, $\mathcal{F}_{k,i}$ is the matched feature point set between map frame $m_{i}$ and user-side frame $c_{k}$, $\mathbf{u}^{m_{i}}_{f}$ and $\mathbf{u}^{c_{k}}_{f}$ are the point coordinates respectively, ${\bm{\Omega}_{p}}$, ${\Omega_{\lambda}}$, ${\bm{\Omega}_{c}}$ are the uncertainties. Notice that the map frame poses are fixed due to unobservability of the reference.
Figure 7: Illustration of how depth uncertainty of the map frame affects the user pose estimation accuracy. In the figure, $\delta s$ denotes the scale uncertainty of the depth map, $\delta p$ denotes the position error of user-side localization. An approximate $\delta p\propto\delta s\cdot|\mathbf{t}^{m_{i}}_{c_{k}}|$ relationship can be found.
@@ VI Experiments
Figure 8: Vehicle trajectories and example images of the experimental dataset. The data sequence collected in 2022/10/23 (5400 sec duration, 30 km mileage, 16 km one-way road coverage) is used for mapping, while the data sequence collected in 2023/04/12 (1800 sec duration, 10 km mileage, 4.9 km one-way road coverage) is used for localization.
@@ VI-A Mapping Performance
For large-scale mapping, obtaining precise global poses is fundamental. Nevertheless, this is a challenging task considering the intermittent degradation of GNSS observations in the experimental scenario. In Fig. 9, we show the pose estimation results of different approaches, including: 1) GNSS RTK, 2) VINS-Fusion (monocular VIO + RTK, final optimized trajectory) and 3) the proposed multi-sensor DBA scheme in real-time mode and global mode respectively. As can be found, the GNSS RTK results are noisy. Although decimeter-level accuracy is intermittently met, gross errors occur frequently. By integrating RTK with VIO, the VINS-Fusion scheme suppresses the gross errors to some extent. For the proposed DBA (real-time) scheme, by integrating visual DBA, inertial information and RTK in a tightly-coupled way, the GNSS errors are better mitigated. Furthermore, with the global optimization presented in Sect. IV-A, the position estimation error the DBA (global) scheme is mostly kept to below 1 m throughout the sequence, which ensures the decimeter-level lower bound of user-side absolute localization.
@@
    3 37.67 MB (JPEG, 60) + 3.41 MB (Global Desc.) + 7.76 MB (Depth)
    4 25.81 MB (JPEG, 60) + 2.32 MB (Global Desc.) + 5.30 MB (Depth)
@@
On the other hand, the generation of map data is looked into. As baselines, we test ORBSLAM3 and HLoc for the mapping task, which are both commonly used pipelines for visual mapping and re-localization. For ORBSLAM3, we simply use the monocular visual-inertial mode to process the sequence and save the map data in the Atlas format. The map storage is over 6 GB for the whole sequence, which is relatively heavy for practical applications. As for HLoc, we found it hard to achieve such large-scale SfM through the default COLMAP[54] + SIFT[55] pipeline. Instead, we directly use the ground-truth camera poses and perform feature extraction, matching and triangulation based on Superpoint and SuperGlue. The map storage is over 30 GB, which is 5 times larger than the ORBSLAM3 map. Such heavy map makes it hard to be applied to large-scale scenarios.
@@ VI-B Localization Performance
As to user-side localization, we check the coarse-to-fine localization performance of the proposed system.
@@ VI-B1 Coarse Localization
We also list the results based on a more sparse map with co-visibility threshold $\Xi=0.3$. Generally, the recall is noticeably lower than the $\Xi=0.4$ case, while the EigenPlaces50 (SSS10) scheme can still achieve 93.78%/100% 10m/20m recall.
@@ VI-B2 Fine Localization
As to the evaluation of fine localization, we test the re-localization modes of ORB-SLAM3 and HLoc as the baselines, and also test GNSS-based schemes for reference, including single point positioning (SPP) and VINS-Fusion (stereo VIO + SPP). For the proposed SF-Loc, we test both single-frame and multi-frame re-localization modes. For multi-frame re-localization, the user-side relative poses are provided by stereo VIO. The factor graph optimization implementation in Sect. V-B is used. For comparison, we also test PnP with RANSAC upon the SF-Loc map.
For ORB-SLAM3 and HLoc, the coarse-to-fine localization is performed based on their own pipeline and map data (corresponding to TABLE I). For SF-Loc, the map data generated with different co-visibility thresholds ($\Xi=0.3/0.4$) are considered, and the coarse localization results of EigenPlaces50 (SSS10) are used.
@@
Finally, we show the time consumption statistics of user-side localization in TABLE IV, which is tested on a laptop with Nvidia RTX 4080 GPU.
TABLE IV: Time Cost of Different Modules in User-Side Localization.
@@ VII Conclusion
In this paper, we propose a vision-centered mapping and localization framework based on the map representation of visual structure frames. The highlight of the system is the lightweight map storage by making the frame data compact and keeping the sparsity, while can still support high-recall, high-accuracy localization through the utilization of user-side multi-frame information.